OpenStudy (anonymous):

The answer is C, but how do you get it? The high school band wants to sell two types of cookies, chocolate chip and peanut butter, as a fundraiser. A dozen chocolate chip cookies requires 2 cups of flour and 1 egg. A dozen peanut butter cookies uses 3 cups of flour and 4 eggs. The club has 90 cups of flour and 80 eggs on hand. The profit on the chocolate chip cookies is $1 per dozen and on the peanut butter is $1.50 per dozen. If they want to offer both types of cookies, how many of each cookie should the club make to maximize profits?

OpenStudy (anonymous):

a. infeasible solutions b. 30 dozen chocolate chip 25 dozen peanut butter c. 24 dozen chocolate chip 14 dozen peanut butter d. 20 dozen chocolate chip 18 dozen peanut butter

OpenStudy (tkhunny):

Name, Quantize, and Translate! C = Number of Dozens Chocolate Chip Cookies (2 flour 1 egg) P = Number of Dozens Peanut Butter Cookies (3 flour 4 eggs) "90 cups of flour" 2C + 3P <= 90 "80 eggs" C + 4P <= 80 "profit on the chocolate chip cookies is $1 per dozen and on the peanut butter is $1.50 per dozen" Profit C(1.00) + P(1.50) You just have to drag through it. There is no perfect way. They way you think about it that you understand is the way you should proceed.

OpenStudy (anonymous):

So you can graph the equations for the eggs and flour 2C + 3P <= 90 C + 4P <= 80 and of their vertices that fall in quadrant 1, you can plug those coordinates into the equation C(1.00) + P(1.50) and that will get you the answer C, right? Tell me if what I typed is confusing.

OpenStudy (tkhunny):

I think you have it. Just show a little more fortitude and read the problem very carefully. There is a LOT of information.